OpenStudy (anonymous):

what is the value of the range of the function f(x)=x^2+2 for the domain value 1/4 WILL MEDAL AND FAN

OpenStudy (anonymous):

a. y = 2 b. y = 2 1/4 c. y = 2 1/8 d. y = 2 1/16

Directrix (directrix):

f(x)=x^2+2 f ( 1/4) = (1/4)^2 + 2 Wherever x is in the function statement, replace that x by 1/4. So, crank out (1/4)^2 + 2 and it should be your answer. @iogenmrfngpg

Directrix (directrix):

I see the answer among your options. Come claim it.